### v1

# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

Abstract
The field of long-term time series forecasting demands handling extensive look-back windows and long-range prediction steps, posing significant challenges for RNN-based methodologies. Among these, SegRNN, a robust RNN-driven model, has gained considerable attention in LTSF analysis for achieving state-of-the-art results while maintaining a remarkably streamlined architecture. Concurrently, the Mamba structure has demonstrated its advantages in small to medium-sized models due to its capability for information selection. This study introduces a variant of SegRNN that preprocesses information using a fine-tuned single-layer Mamba structure. Additionally, it incorporates implicit segmentation and residual structures into the model’s encoding section to further reduce the inherent data iterative cycles of RNN architectures and implicitly integrate inter-channel correlations. This variant, named MSegRNN, utilizes the Mamba structure to select useful information, resulting in a transformed sequence. The linear-strategy-adapted derivative retains the superior memory efficiency of the original SegRNN while demonstrating enhanced performance. Empirical evaluations on real-world LTSF datasets demonstrate the superior performance of our model, thereby contributing to the advancement of LTSF methodologies.
Keywords Long Time Series Analysis  $\cdot$ RNN  $\cdot$ Mamba  $\cdot$ Implicit Segment
1 Introduction
Long-term series forecasting is the process of predicting the future values of a sequence over an extended period. This task primarily aims to forecast long-term trends, which hold significant importance in domains such as climate prediction (Aghelpour et al., 2019), decision support (Nooruldeen et al., 2022), and policy formulation (Zhang et al., 2022). Long-term sequence data often exhibit higher levels of uncertainty, which can result in reduced prediction accuracy. Additionally, longer forecasting horizons require models to incorporate a broader historical context for accurate predictions, thereby increasing the complexity of modeling. Recent advancesments in long-term sequence analysis have shifted towards deep learning methods, highlighting the crucial need to develop diverse and effective time-series methodologies.
Long-term time series analysis and forecasting comprise two critical components: (1) trend identification and analysis, which involves recognizing long-term directional changes in the data, This component is crucial for understanding the behavior of time series, enabling the comprehension of past patterns and the prediction of potential future developments. (2) Uncertainty and noise management: Due to the extended time range of long-term time series data, they often exhibit increased levels of uncertainty and noise. Effectively managing and mitigating these uncertainties and noises is vital for improving the accuracy of predictions.
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Nie et al., 2022; Wu et al., 2021; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
The Mamba architecture, proposed by Gu and Dao(Gu and Dao, 2023), constitutes a state space model tailored for efficient sequence modeling. It integrates parameterized matrices and a hardware-aware parallel computing algorithm, achieving impressive results in tasks such as language modeling and time series forecasting. Mamba is notable for its ability to capture long-term dependencies while maintaining linear computational complexity, making it a strong competitor to Transformer-based models. This capability is facilitated through its adept management of state space models and the implementation of a selective state space mechanism designed to optimize performance.
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series analysis tasks. SegRNN employ a channel-independent strategy to enhance its effectiveness. Nevertheless, this strategy fails to effectively handle noise and capture the correlations between variables. Additionally, even with the adoption of the segmentation strategy, SegRNN still faces the issue of information loss within the recurrent structure.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segment strategy. Specifically, we propose a long-term time series prediction model based on Mamba transformation and linear encoding, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
  • Introducing the Mamba structure to effectively extract the information
  • Utilizes information between channels to promote the infomation
  • The implementation of residual network in the encoding phase reduces the accumulation of errors during encoding stage.
The structure of this paper is organized as follow: Section 2 formally presents our problem statement. Section 3 discusses related work. Section 4 elaborates on our method of introducing Mamba structure and other technologys. Section 5 describes our experiments and results. Section 6 concludes the paper and suggests future research directions.
2 Problem statement
The analysis of long-term time series involves addressing the challenge of predicting future values of multivariate time series based on historical data. Typically, the problem of time series analysis can be formulated as follows:
$Y=F(X)+\xi$ (1)
In which $X\in\mathbb{R}^{L\times C}$ represents multi-variable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbb{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and H indicates the time step to be predicted. Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
3 Related Work
A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. In this section, we introduce literature relevant to our work. These methods can be broadly divided into the following four categories:
3.1 Transformer Models
Transformer models, originally designed for sequence-to-sequence tasks, have achieved significant achievements accross various domains. The core of applying Transformer to long-time series tasks is to use the attention mechanism to capture the long-term dependencies (Hao and Cao, 2020). Extensive research has been devoted to advancing the application of the Transformer architecture in long-time series analysis. Autoformer (Wu et al., 2021) decomposes the sequence into basic blocks within the model, thus gradually decomposing complex time series and enhancing learning ability for complex patterns. Informer (Zhou et al., 2021) constructs a ProbSparse self-attention mechanism, adopting self-attention distillation and a generative decoder to address the high time complexity, large memory consumption, and encoder architecture limitations faced in long-time series analysis. Crossformer (Zhang and Yan, 2022) embeds time series into two-dimensional vector through the Dimension-segment-wise structure, effectively captures dependencies across time and dimensions through the Two-stage-attention structure. The segmentation technique used in Crossformer method is considered beneficial for long time series foresting.
3.2 MLP Models
Extensive research has been dedicated to the application of MLP methods in long-term time series forecasting. Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables. These variables are then mapped to the prediction results via linear layers. The Nlinear (Zeng et al., 2023) method adopts a regularization strategy in time series forecasting tasks to enhance performance. Specifically, it subtracts the last value before inputting to the model, mapping through a linear layer, then adding it back for the final prediction. The TiDE (Nie et al., 2022) method introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. Long-term time series analysis methods based on MLP have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus advancing the development of MLP in long-term time series forecasting.
3.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. The Disjoint-CNN (Foumani et al., 2021) approach, while maintaining computational efficiency, enhances prediction accuracy by decomposing one-dimensional convolutional kernels into disjoint temporal and spatial components. Nonetheless, CNN-based methodologies encounter challenges as the forecasting horizon expands. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
3.4 RNN Models
The RNN architecture was initially designed for processing sequence tasks and excel in short-term time series prediction. However, RNN-based methods face error accumulation issues in long-term time series tasks. SegRNN (Lin et al., 2023) aims to mitigate the accumulation of errors in the recurrent structure while preserving sequence information. Specifically, in the encoding stage, the SegRNN method segments the data before encoding. In the decoding stage, it employs a PMF approach for parallel decoding, significantly enhancing prediction accuracy and the spatiotemporal efficiency.
Nonetheless, the SegRNN method is not without its limitations: (1) SegRNN couldn’t effectivly preporcess the data and extract the inforamtion, which may lead to a rapid accumulation of noise. (2)SegRNN adopts fixed segmentation of the original time series, which may result in partial information loss due to its rigidity. (3) SegRNN aims to minimize the number of recurrent iterations while striving to retain sequential information. However, there are still limitations in minimizing the number of recurrent iterations, which restrict the model’s capability. In the following section, we will propose some strategies to address these issues.
4 Method
To address the challenges faced by current methods, we incorporate mamba structure, implicit segmentation and residual networks into the SegRNN framework. This approach (1) first extracts the original data first to contrain noise, (2)segments the original data through linear mapping to reduce information loss caused by fixed segmentation, and (3) introduces residual structures at the encoding layer to mitigate information loss within the recurrent structure.
We first discuss the overall sturcture that we used in Section 4.1, then we discuss the comparison between explicit segmentation and implicit segmentation in Section 4.2, followed by a discussion on how to introduce residual structures into the model during the encoding process of time series data in Section 4.3.
4.1 Model Structure
The Mamba architecture achieves long-term sequence memory and model selection capabilities by setting the SSM structure parameters as trainable parameters dependent on the sequence content. We have adapted the Mamba architecture to address time series preprocessing problems. The basic structure of our model is shown in Figure 1:
Figure 1: Overall structure of MSegRNN model
Several modifications were applied to the Mamba block and SegRNN model to improve performance and flexibility. Specifically, the convolutional layer in Mamba was removed to enhance computational efficiency, with local information being handled by the SegRNN structure. The SegRNN model’s fixed segmentation was replaced with Implicit Segmentation for adaptability, and residual connections were integrated to strengthen the linkage between segmentation and encoding outputs. Further details on these adjustments are discussed in Section 4.2 and Section 4.3.
4.2 Implicit Segmentation
The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series. Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
$n=\frac{L}{w}$ (2)
Subsequently, the segmented model projects through a linear layer:
$\tilde{X}_{w}^{(i)}=W_{prj}\cdot X_{w}^{(i)}$ (3)
where $W_{prj}\in\mathbb{R}^{w\times d}$ denotes the projection matrix of linear mapping, and $d$ represents the dimensionality of the hidden layer within the RNN structure. Undergoing linear projection transformation, the dimensionality of $\tilde{X}_{w}^{(i)}$ is defined as $\mathbb{R}^{C\times n\times d}$.
However, this segmentation strategy has certain limitations. Once the segment length is chosen, the cut-off points are also determined. This may disrupt the continuity of information in time series analysis, thus limiting the model’s capabilities. The implicit segmentation method aims to break this limitation. Specifically, given a time series analysis vector $X^{(i)}\in\mathbb{R}^{L\times C}$, we first expand its last dimension and then swap the first and second dimensions to obtain $\tilde{X}^{(i)}\in\mathbb{R}^{C\times L\times 1}$ for subsequent processing. Then we utlize a linear layer to transform the last dimesntion of $\tilde{X}^{(i)}$ to get $\bar{X}(i)\in\mathbb{R}^{C\times L\times n}$ to represent the segmented matrix, subsequently, we interchange the second and third dimension, still denoted as $\bar{X}(i)$. Applying linear layer to the last dimension of $\bar{X}(i)$ yields $\bar{X}(i)\in\mathbb{R}^{C\times n\times d}$, which is consistent in dimension with the vector obtained via SegRNN methodology.
Notably, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
4.3 Residual Structure for SegRNN
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbb{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
$\mathbf{h}_{0}$ $=\mathbf{0}$ (4)
$\mathbf{h}_{1}$ $=f(\mathbf{h}_{0},\tilde{\mathbf{X}}_{w}^{(i)}[:,1,:];\mathbf{W})$
$\mathbf{h}_{2}$ $=f(\mathbf{h}_{1},\tilde{\mathbf{X}}_{w}^{(i)}[:,2,:];\mathbf{W})$
$\vdots$
$\mathbf{h}_{n}$ $=f(\mathbf{h}_{n-1},\tilde{\mathbf{X}}_{w}^{(i)}[:,n,:];\mathbf{W})$
where $\mathbf{h}_{t}\in\mathbb{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps. This unavoidably limits the model’s ability due to the loss of information during iterations. Therefore, we introduce a residual structure in the model. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$. Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$. Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
Introducing the residual structure allows some information to bypass the recurrent structure and be directly mapped to the encoder output. This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.
5 Experiments
Our experiments are structured as follows. In Section 5.1, we introduce the datasets used for long-term time series analysis. Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison. Section 5.3 presents and analyzes the performance metrics. In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model. Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model. All experiments in this section were conducted on a single NVIDIA 4090 GPU.
5.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, and traffic, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
Table 1: Dataset Basic Attributes
Dataset Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
Time-Series 321 21 7 7 7 7
Frequency 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
Time-Points 26,304 52,696 17,420 17,420 69,680 69,680
5.2 Experimental setup and Baselines
The unified configuration of the model is substantially aligned with the SegRNN approach. The look-back window is set to be 96, implicit Segmentation length is set to be 24 or 48 based on the size of datasets. A single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing. The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30. Dropout rate, learning rate, and batch size vary with the data and the scale of the model.
As baselines, we have selected state-of-the-art and representative models in the long-term time series forecasting domain, comprising the following categories: (i) RNNs: SegRNN; (ii) MLPs: Dlinear; (iii) CNNs: TimesNet; (iv) Transformers: iTransformer, Crossformer, PatchTST.
5.3 Main Result
We selected two evaluation metrics, MSE and MAE, with prediction steps of 96, 192, 336, and 720 time steps. The resulting forecasts are presented in Table 2. Note that all data, except for our method, originate from official code repositories and original papers.
Our method achieved top positions in 37 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics. The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively. This demonstrates the powerful capabilities of our model.
Table 2: Performance metrics for Long-time series analysis.
Datasets Models RNNs Transformers MLPs CNNs
MSegRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.368 0.383 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
192 0.422 0.414 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
336 0.471 0.436 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
720 0.483 0.459 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
ETTh2 96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
192 0.354 0.378 0.359 0.389 0.380 0.400 0.877 0.656 0.388 0.400 0.477 0.476 0.402 0.414
336 0.404 0.416 0.421 0.436 0.428 0.432 1.043 0.731 0.426 0.433 0.594 0.541 0.452 0.452
720 0.407 0.428 0.432 0.455 0.427 0.445 1.104 0.763 0.431 0.446 0.831 0.657 0.462 0.468
ETTm1 96 0.314 0.345 0.330 0.369 0.334 0.368 0.404 0.426 0.329 0.367 0.345 0.372 0.338 0.375
192 0.361 0.374 0.369 0.392 0.377 0.391 0.450 0.451 0.367 0.385 0.380 0.389 0.374 0.387
336 0.388 0.392 0.399 0.412 0.426 0.420 0.532 0.515 0.399 0.410 0.413 0.413 0.410 0.411
720 0.448 0.430 0.454 0.443 0.491 0.459 0.666 0.589 0.454 0.439 0.474 0.453 0.478 0.450
ETTm2 96 0.170 0.248 0.173 0.255 0.180 0.264 0.287 0.366 0.175 0.259 0.193 0.292 0.187 0.267
192 0.233 0.292 0.237 0.298 0.250 0.309 0.414 0.492 0.241 0.302 0.284 0.362 0.249 0.309
336 0.294 0.331 0.296 0.336 0.311 0.348 0.597 0.542 0.305 0.343 0.369 0.427 0.321 0.351
720 0.391 0.389 0.389 0.407 0.412 0.407 1.730 1.042 0.402 0.400 0.554 0.522 0.408 0.403
Electricity 96 0.149 0.242 0.151 0.245 0.148 0.240 0.219 0.314 0.181 0.270 0.197 0.282 0.168 0.272
192 0.165 0.259 0.164 0.258 0.162 0.253 0.231 0.322 0.188 0.274 0.196 0.285 0.184 0.289
336 0.181 0.275 0.180 0.277 0.178 0.269 0.246 0.337 0.204 0.293 0.209 0.301 0.198 0.300
720 0.218 0.307 0.218 0.313 0.225 0.317 0.280 0.363 0.246 0.324 0.245 0.333 0.220 0.320
Weather 96 0.155 0.193 0.165 0.227 0.174 0.214 0.158 0.230 0.177 0.210 0.196 0.255 0.172 0.220
192 0.206 0.244 0.211 0.273 0.221 0.254 0.206 0.277 0.225 0.250 0.237 0.296 0.219 0.261
336 0.260 0.288 0.270 0.318 0.278 0.296 0.272 0.335 0.278 0.290 0.283 0.335 0.280 0.306
720 0.341 0.340 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
Count 37 7 6 1 1 0 0
5.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. Using the weather dataset as an example, we performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
Table 3: Ablation study on different structure
Datasets Models M&LR&SegRNN LR&SegRNN M&SegRNN SegRNN
MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.155 0.193 0.163 0.202 0.191 0.232 0.165 0.227
192 0.206 0.244 0.207 0.242 0.243 0.277 0.211 0.273
336 0.260 0.288 0.264 0.287 0.305 0.324 0.270 0.318
720 0.341 0.340 0.350 0.342 0.380 0.377 0.357 0.376
Average 0.241 0.266 0.246 0.268 0.280 0.303 0.251 0.298
The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall. This is followed by the model that includes only implicit segmentation and residual structures. Interestingly, merely adding the Mamba structure does not significantly enhance model performance. This may be due to a synergistic effect between the two structures. Consequently, the ablation study confirms the superiority of our proposed structure.
5.5 Model efficiency
Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Figure 2.
Figure 2: Mean memory consumpation in weather datasets for MSegRNN and SegRNN
As shown in the figure, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
6 Conclusion
In this work, we present the MSegRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of MSegRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the MSegRNN method excels with corresponding empirical research.

### v2

@@
 4.2 Implicit Segmentation
-The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series.
+The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, As shown in Figure 2, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series.
 Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
@@
 Applying linear layer to the last dimension of $\bar{X}(i)$ yields $\bar{X}(i)\in\mathbb{R}^{C\times n\times d}$, which is consistent in dimension with the vector obtained via SegRNN methodology.
+Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation
 Notably, our proposed method does not explicitly define the segmentation length.
 Instead, it achieves data segmentation and dimensionality transformation through dual linear projections.
-The incorporation of linear fully connected layers facilitates a more continuous transformation of information.
+The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information.
 It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
@@
 This unavoidably limits the model’s ability due to the loss of information during iterations.
-Therefore, we introduce a residual structure in the model.
+Therefore, we introduce a residual structure in the model, as shown in Figure 3.
 For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$.
@@
 Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
+Figure 3: Residual block for SegRNN
 Introducing the residual structure allows some information to bypass the recurrent structure and be directly mapped to the encoder output.
@@
 Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model.
-All experiments in this section were conducted on a single NVIDIA 4090 GPU.
+All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
 5.1 Datasets
@@
 The unified configuration of the model is substantially aligned with the SegRNN approach.
-The look-back window is set to be 96, implicit Segmentation length is set to be 24 or 48 based on the size of datasets.
-A single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing.
+The look-back window is set to be 96, a single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing.
 The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30.
@@
 Note that all data, except for our method, originate from official code repositories and original papers.
-Our method achieved top positions in 37 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics.
+Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics.
 The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively.
@@
 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 96 0.368 0.383 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
-192 0.422 0.414 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
-336 0.471 0.436 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
-720 0.483 0.459 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
+ETTh1 96 0.365 0.384 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
+192 0.415 0.413 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
+336 0.463 0.441 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
+720 0.468 0.460 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
 ETTh2 96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
@@
 720 0.341 0.340 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
-Count 37 7 6 1 1 0 0
+Count 36 7 6 1 1 0 0
+We subsequently conducted experiments to observe how the model’s capability varies with changes in the look-back window.
+The hyperparameter settings of the method were kept consistent with look-back window 96, and the horizon was uniformly set to 336.
+The evaluation was performed on the ETTh2 dataset, with the MSE and MAE losses of the method shown in the Figure 4:
+Figure 4: Comparsion of SegRNN and MSegRNN with vary look-back window length
+Our model exhibited smaller losses across various look-back windows, demonstrating the superiority of our method.
+Notably, the advantage of our method is more pronounced with smaller look-back windows.
+As the look-back window increases, this advantage diminishes, possibly due to noise introduced by segmentation affecting the model.
+Nevertheless, our model maintained a considerable advantage.
 5.4 Ablation study
@@
 336 0.260 0.288 0.264 0.287 0.305 0.324 0.270 0.318
-720 0.341 0.340 0.350 0.342 0.380 0.377 0.357 0.376
-Average 0.241 0.266 0.246 0.268 0.280 0.303 0.251 0.298
+720 0.341 0.335 0.350 0.342 0.380 0.377 0.357 0.376
+ETTh1 96 0.365 0.384 0.377 0.387 0.480 0.446 0.368 0.395
+192 0.415 0.413 0.423 0.416 0.520 0.472 0.408 0.419
+336 0.463 0.441 0.475 0.437 0.543 0.491 0.444 0.440
+720 0.468 0.460 0.476 0.454 0.549 0.515 0.446 0.457
+ETTh2 96 0.275 0.326 0.281 0.327 0.326 0.368 0.278 0.335
+192 0.354 0.378 0.368 0.385 0.418 0.420 0.359 0.389
+336 0.404 0.416 0.402 0.416 0.462 0.456 0.421 0.436
+720 0.407 0.428 0.421 0.435 0.467 0.472 0.432 0.455
+ETTm1 96 0.314 0.345 0.315 0.346 0.535 0.454 0.330 0.369
+192 0.361 0.374 0.362 0.376 0.556 0.469 0.369 0.392
+336 0.388 0.392 0.390 0.395 0.576 0.480 0.399 0.412
+720 0.448 0.430 0.446 0.427 0.631 0.509 0.454 0.443
+ETTm2 96 0.170 0.248 0.173 0.250 0.215 0.298 0.173 0.255
+192 0.233 0.292 0.235 0.293 0.283 0.339 0.237 0.298
+336 0.294 0.331 0.297 0.332 0.352 0.381 0.296 0.336
+720 0.391 0.389 0.392 0.388 0.463 0.441 0.389 0.407
+Average 0.335 0.355 0.341 0.357 0.425 0.411 0.340 0.372
 The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall.
@@
 The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments.
-Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Figure 2.
-Figure 2: Mean memory consumpation in weather datasets for MSegRNN and SegRNN
-As shown in the figure, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
+Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Table 4.
+Table 4: Memory Usage of Two Methods
+Method Memory Usage (GB)
+96 192 336 720
+MSegRNN 0.98 1.17 1.56 2.48
+SegRNN 0.91 1.14 1.53 2.41
+Compare +7.7% +2.6% +2.0% +2.9%
+As shown in the table, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
 6 Conclusion
@@
 Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the MSegRNN method excels with corresponding empirical research.
+Appendix
+Appendix A Datasets
+For our Long-term Time Series Forecasting (LTSF) study, we employed several widely-used multivariate datasets, including:
+ETTs: This collection consists of electricity transformer data from two counties in China, spanning two years.
+The datasets include various subsets: ETTh1 and ETTh2 provide hourly data, and ETTm1 and ETTm2 offer 15-minute interval data.
+Each entry encompasses the target variable "oil temperature" alongside six power load indicators.
+Electricity: This dataset contains hourly electricity consumption records for 321 clients from 2012 to 2014, offering insights into consumer energy usage patterns over time.
+Weather: This dataset captures 21 meteorological variables, such as temperature and humidity, at 10-minute intervals throughout 2020.
+It serves as a comprehensive resource for weather prediction and climate studies.
+These datasets are instrumental for benchmarking and advancing forecasting models in the context of long-term time series analysis.
+Appendix B Code Framework
+The implementation of our model framework is primarily derived from the official SegRNN code and the minimal implementation of Mamba.
+This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output.
+We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
+The official SegRNN code is available at:https://github.com/lss-1138/SegRNN.
+The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal
+Appendix C Detail Configuration
+The main hyperparameter settings of our model are shown in Table 5.
+The parameter $d_{\text{state}}$ is a hyperparameter in the Mamba architecture, which is detailed in Mamba [Gu and Dao, 2023].
+The parameter settings and results for other methods are taken from their original papers or official code repositories.
+Table 5: Detail configuration of MSegRNN
+Datasets d_state seg_len d_model learning_rate dropout batch_size
+ETTh1 2 12 512 0.0003 0.1 256
+ETTh2 4 24 512 0.0015 0.5 256
+ETTm1 4 24 512 0.0010 0.1 256
+ETTm2 4 24 512 0.0010 0.1 256
+Weather 4 24 512 0.0007 0.1 64
+Electricity 2 48 512 0.0012 0.01 128
+We also employed a different learning rate adjustment strategy.
+The learning rate remained constant for the first 15 epochs and then gradually multiplied by 0.9.
+This approach actually increases the learning rate compared to the adjustment method used in SegRNN.
+Appendix D Comparison with extra methods
+We observed that some methods also utilized Mamba for information augmentation, we selected several long-term time series analysis models based on the Mamba architecture: Time-Machine[Ahamed and Cheng, 2024], Bi-Mamba4TS[Liang et al., 2024], and DT-Mamba[Wu et al., 2024], for comparison.
+The results are from official paper, respectivly .
+The result is shown in Table 6.
+Table 6: Perfromance of different models
+Datasets Models MSegRNN TimeMachine Bi-Mamba4TS DTMamba
+MSE MAE MSE MAE MSE MAE MSE MAE
+Weather 96 0.155 0.193 0.164 0.208 0.164 0.212 0.171 0.218
+192 0.206 0.244 0.211 0.250 0.214 0.256 0.220 0.257
+336 0.260 0.288 0.256 0.290 0.269 0.296 0.274 0.296
+720 0.341 0.340 0.342 0.343 0.348 0.349 0.349 0.346
+ETTh1 96 0.365 0.384 0.364 0.387 0.376 0.403 0.386 0.399
+192 0.415 0.414 0.415 0.416 0.411 0.425 0.426 0.424
+336 0.463 0.441 0.429 0.421 0.455 0.445 0.480 0.450
+720 0.468 0.460 0.458 0.453 0.460 0.464 0.484 0.470
+ETTh2 96 0.275 0.326 0.275 0.334 0.289 0.341 0.290 0.340
+192 0.354 0.378 0.349 0.381 0.367 0.389 0.366 0.392
+336 0.404 0.416 0.340 0.381 0.410 0.424 0.380 0.409
+720 0.407 0.428 0.411 0.433 0.421 0.439 0.416 0.437
+ETTm1 96 0.314 0.345 0.317 0.355 0.312 0.354 0.325 0.360
+192 0.361 0.374 0.357 0.378 0.358 0.383 0.375 0.386
+336 0.388 0.392 0.379 0.399 0.388 0.404 0.396 0.405
+720 0.448 0.430 0.445 0.436 0.444 0.433 0.454 0.442
+ETTm2 96 0.170 0.248 0.175 0.256 0.174 0.259 0.177 0.259
+192 0.233 0.292 0.239 0.299 0.240 0.304 0.240 0.300
+336 0.294 0.331 0.287 0.332 0.303 0.345 0.310 0.345
+720 0.391 0.389 0.371 0.385 0.402 0.407 0.395 0.394
+Electricity 96 0.149 0.242 0.142 0.236 0.142 0.238 0.166 0.256
+192 0.165 0.259 0.158 0.250 0.157 0.253 0.178 0.268
+336 0.181 0.275 0.172 0.268 0.172 0.271 0.197 0.289
+720 0.218 0.307 0.207 0.298 0.200 0.297 0.243 0.326
+It can be observed that our method performs notably well on the weather, ETTh2, and ETTm2 datasets, while the TimeMachine method shows significant advantages on the ETTh1 and Electricity datasets.
+Additionally, the Bi-Mamba4TS method also demonstrates a clear advantage on the Electricity dataset.

### v3

@@
 Abstract
-The field of long-term time series forecasting demands handling extensive look-back windows and long-range prediction steps, posing significant challenges for RNN-based methodologies.
-Among these, SegRNN, a robust RNN-driven model, has gained considerable attention in LTSF analysis for achieving state-of-the-art results while maintaining a remarkably streamlined architecture.
-Concurrently, the Mamba structure has demonstrated its advantages in small to medium-sized models due to its capability for information selection.
-This study introduces a variant of SegRNN that preprocesses information using a fine-tuned single-layer Mamba structure.
-Additionally, it incorporates implicit segmentation and residual structures into the model’s encoding section to further reduce the inherent data iterative cycles of RNN architectures and implicitly integrate inter-channel correlations.
-This variant, named MSegRNN, utilizes the Mamba structure to select useful information, resulting in a transformed sequence.
-The linear-strategy-adapted derivative retains the superior memory efficiency of the original SegRNN while demonstrating enhanced performance.
-Empirical evaluations on real-world LTSF datasets demonstrate the superior performance of our model, thereby contributing to the advancement of LTSF methodologies.
+Long time series forecasting aims to utilize historical information to forecast future states over extended horizons.
+Traditional RNN-based series forecasting methods struggle to effectively address long-term dependencies and gradient issues in long time series problems.
+Recently, SegRNN has emerged as a leading RNN-based model tailored for long-term series forecasting, demonstrating state-of-the-art performance while maintaining a streamlined architecture through innovative segmentation and parallel decoding techniques.
+Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure.
+To address these issues, we propose the MSegRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase.
+Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure.
+To extract information more effectively, we further integrate the Mamba architecture to enhance time series information extraction.
+Experiments on several real-world long time series forecasting datasets demonstrate that our model surpasses the performance of current state-of-the-art models.
 Keywords Long Time Series Analysis  $\cdot$ RNN  $\cdot$ Mamba  $\cdot$ Implicit Segment
@@
 In the context of long-term time series analysis, it is essential to capture the sequential relationships between data.
-Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Nie et al., 2022; Wu et al., 2021; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023).
+Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023a; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Das et al., 2023b; Nie et al., 2022; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023).
 CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation.
@@
 (Lin et al., 2023) proposed the segmentation-based SegRNN model.
-This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series analysis tasks.
-SegRNN employ a channel-independent strategy to enhance its effectiveness.
-Nevertheless, this strategy fails to effectively handle noise and capture the correlations between variables.
-Additionally, even with the adoption of the segmentation strategy, SegRNN still faces the issue of information loss within the recurrent structure.
-In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segment strategy.
-Specifically, we propose a long-term time series prediction model based on Mamba transformation and linear encoding, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
-  • Introducing the Mamba structure to effectively extract the information
-  • Utilizes information between channels to promote the infomation
+This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series foreasting tasks.
+SegRNN employs a segmented approach to time series, which may disrupt the continuity of the sequences.
+Although the segmentation strategy effectively balances sequence information and iteration count, information loss within the recurrent structure occurs as the prediction horizon extends, thereby limiting the model’s capability.
+In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy.
+Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
+  • Introducing the Mamba structure to effectively preprocessing the information.
+  • Utilizing implicit segmentation for denser information exchange during the segmentation phase.
   • The implementation of residual network in the encoding phase reduces the accumulation of errors during encoding stage.
-The structure of this paper is organized as follow: Section 2 formally presents our problem statement.
-Section 3 discusses related work.
-Section 4 elaborates on our method of introducing Mamba structure and other technologys.
-Section 5 describes our experiments and results.
-Section 6 concludes the paper and suggests future research directions.
 2 Problem Statement
@@
 $Y=F(X)+\xi$ (1)
-In which $X\in\mathbb{R}^{L\times C}$ represents multi-variable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbb{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and H indicates the time step to be predicted.
+In which $X\in\mathbf{R}^{L\times C}$ represents multivariable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbf{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and $H$ indicates the time step to be predicted.
 Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
-3 Related Work
-A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis.
-In this section, we introduce literature relevant to our work.
-These methods can be broadly divided into the following four categories:
+A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. These methods can be broadly divided into the following categories:
 2.1 Transformer Models
-Transformer models, originally designed for sequence-to-sequence tasks, have achieved significant achievements accross various domains.
-The core of applying Transformer to long-time series tasks is to use the attention mechanism to capture the long-term dependencies (Hao and Cao, 2020).
-Extensive research has been devoted to advancing the application of the Transformer architecture in long-time series analysis.
-Autoformer (Wu et al., 2021) decomposes the sequence into basic blocks within the model, thus gradually decomposing complex time series and enhancing learning ability for complex patterns.
-Informer (Zhou et al., 2021) constructs a ProbSparse self-attention mechanism, adopting self-attention distillation and a generative decoder to address the high time complexity, large memory consumption, and encoder architecture limitations faced in long-time series analysis.
-Crossformer (Zhang and Yan, 2022) embeds time series into two-dimensional vector through the Dimension-segment-wise structure, effectively captures dependencies across time and dimensions through the Two-stage-attention structure.
-The segmentation technique used in Crossformer method is considered beneficial for long time series foresting.
+The core of applying Transformer models to long-term time series tasks lies in their ability to capture long-term dependencies using the attention mechanism (Hao and Cao, 2020).
+The general form of the self-attention mechanism is given by:
+$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (2)
+Significant research efforts have advanced the application of Transformer architectures in long-term time series analysis.
+For instance, PatchTST(Nie et al., 2022) introduces a method of segmenting time series into patches, reducing computational complexity and improving the model’s ability to capture local semantic information.
+It also employs self-supervised learning, masking random patches and training the model to reconstruct them, enhancing forecasting accuracy.
+iTransformer(Liu et al., 2023) addresses scalability by integrating hierarchical attention mechanisms and temporal convolutional networks, capturing both short-term and long-term dependencies more effectively.
+This lightweight architecture balances accuracy and efficiency, making it suitable for real-world applications.
+Crossformer (Zhang and Yan, 2022) employs a Dimension-segment-wise structure to embed time series into a two-dimensional vector, effectively capturing dependencies across time and dimensions through its Two-stage-attention structure.
+The segmentation technique utilized in Crossformer is particularly advantageous for long-term time series forecasting.
 2.2 MLP Models
-Extensive research has been dedicated to the application of MLP methods in long-term time series forecasting.
+MLP methods have garnered significant attention in long-term time series forecasting due to their robust performance.
 Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables.
 These variables are then mapped to the prediction results via linear layers.
-The Nlinear (Zeng et al., 2023) method adopts a regularization strategy in time series forecasting tasks to enhance performance.
-Specifically, it subtracts the last value before inputting to the model, mapping through a linear layer, then adding it back for the final prediction.
-The TiDE (Nie et al., 2022) method introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results.
-Long-term time series analysis methods based on MLP have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus advancing the development of MLP in long-term time series forecasting.
+The Dlinear model can be described as:
+$\hat{Y}=W_{1}(MA(X))+W_{2}(R(X))$ (3)
+where $MA(X)$ represents the trend component obtained by the moving average kernel, $R(X)$ represents the remainder component, and $W_{1}$, $W_{2}$ are the weights of the linear layers.
+TiDE (Das et al., 2023b) introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results.
+The core operation of TiDE involves projecting features and encoding them through residual blocks, which can be described as:
+$\hat{Y}=\text{ResBlock}(\text{ReLU}(W_{f}X+b_{f})+\text{Covariates})$ (4)
+where $W_{f}$ and $b_{f}$ are the weights and biases of the projection layer, respectively.
+The ResBlock operation encapsulates the residual connections used to maintain the integrity of the encoded features while integrating the covariates.
+MLP-based methods for long-term time series analysis have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus significantly advancing the development of MLP in long-term time series forecasting.
 2.3 CNN Models
 The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features.
-The Disjoint-CNN (Foumani et al., 2021) approach, while maintaining computational efficiency, enhances prediction accuracy by decomposing one-dimensional convolutional kernels into disjoint temporal and spatial components.
-Nonetheless, CNN-based methodologies encounter challenges as the forecasting horizon expands.
 MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series.
@@
 These methodologies have attained remarkable results in the domain of long-term time series forecasting.
-3.4 RNN Models
-The RNN architecture was initially designed for processing sequence tasks and excel in short-term time series prediction.
-However, RNN-based methods face error accumulation issues in long-term time series tasks.
-SegRNN (Lin et al., 2023) aims to mitigate the accumulation of errors in the recurrent structure while preserving sequence information.
-Specifically, in the encoding stage, the SegRNN method segments the data before encoding.
-In the decoding stage, it employs a PMF approach for parallel decoding, significantly enhancing prediction accuracy and the spatiotemporal efficiency.
-Nonetheless, the SegRNN method is not without its limitations: (1) SegRNN couldn’t effectivly preporcess the data and extract the inforamtion, which may lead to a rapid accumulation of noise.
-(2)SegRNN adopts fixed segmentation of the original time series, which may result in partial information loss due to its rigidity.
-(3) SegRNN aims to minimize the number of recurrent iterations while striving to retain sequential information.
-However, there are still limitations in minimizing the number of recurrent iterations, which restrict the model’s capability.
-In the following section, we will propose some strategies to address these issues.
+2.4 SegRNN Model
+As the only RNN-based model specifically designed for long-term time series tasks, SegRNN has achieved remarkable performance.
+SegRNN addresses the challenge of retaining sequence information while minimizing information loss within the recurrent structure through two key methods:
+  • SegRNN replaces point-wise iterations with segment-wise iterations.
+This approach balances the trade-off between sequence information retention and the number of recurrent iterations, leading to improved outcomes.
+  • SegRNN replaces recurrent multi-step forecasting with parallel multi-step forecasting. This parallel decoding approach enhances the model’s predictive capability and inference speed.
+By combining these two strategies, SegRNN achieves excellent performance within a lightweight model structure, along with superior inference speed.
 3 Method
-To address the challenges faced by current methods, we incorporate mamba structure, implicit segmentation and residual networks into the SegRNN framework.
-This approach (1) first extracts the original data first to contrain noise, (2)segments the original data through linear mapping to reduce information loss caused by fixed segmentation, and (3) introduces residual structures at the encoding layer to mitigate information loss within the recurrent structure.
-We first discuss the overall sturcture that we used in Section 4.1, then we discuss the comparison between explicit segmentation and implicit segmentation in Section 4.2, followed by a discussion on how to introduce residual structures into the model during the encoding process of time series data in Section 4.3.
-4.1 Model Structure
-The Mamba architecture achieves long-term sequence memory and model selection capabilities by setting the SSM structure parameters as trainable parameters dependent on the sequence content.
-We have adapted the Mamba architecture to address time series preprocessing problems.
-The basic structure of our model is shown in Figure 1:
+3.1 Overview
+Previous work, such as SegRNN, has achieved remarkable results by employing segmentation and parallel decoding techniques.
+However, the fixed segmentation approach restricts the model’s capacity and disrupts data continuity.
+Additionally, segmentation methods still face the challenge of cumulative information processing within the recurrent structure.
+In this paper, we propose the MSegRNN method to overcome these limitations through three strategies, as illustrated in Figure 1.
+We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage.
+Subsequently, we incorporate a residual structure to enable partial information bypassing the recurrent structure, thereby reducing information loss in recurrent structure.
+We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
 Figure 1: Overall structure of MSegRNN model
-Several modifications were applied to the Mamba block and SegRNN model to improve performance and flexibility.
-Specifically, the convolutional layer in Mamba was removed to enhance computational efficiency, with local information being handled by the SegRNN structure.
-The SegRNN model’s fixed segmentation was replaced with Implicit Segmentation for adaptability, and residual connections were integrated to strengthen the linkage between segmentation and encoding outputs.
-Further details on these adjustments are discussed in Section 4.2 and Section 4.3.
+3.2 Mamba structure
+The Mamba architecture, initially proposed by Gu and Dao (2023), has demonstrated outstanding performance in various fields such as large language models and long time series forecasting.
+By setting model parameters to be content-related, the Mamba architecture achieves powerful data selection capabilities, which are crucial for time series forecasting tasks.
+Therefore, we utilize the Mamba architecture at the initial stage to preprocess and enhance our time series data.
+It is important to note that our model does not incorporate the convolutional layer of the Mamba architecture except for weather dataset.
+This difference is mainly based on performance considerations, as we believe that local information can be effectively captured through segmentation.
 3.3 Implicit Segmentation
-The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, As shown in Figure 2, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series.
-Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
+The segmentation method employed by SegRNN, which we refer to as explicit segmentation to distinguish it from our segmentation approach, is implemented by truncating the time dimension.
+For a given time series $X^{(i)}\in\mathbf{R}^{L\times C}$, selecting a segmentation length $w$ results in segmented vectors of size $X_{w}^{(i)}\in\mathbf{R}^{n\times w\times C}$, where $n$ represents the number of segments.
 $n=\frac{L}{w}$ (5)
-Subsequently, the segmented model projects through a linear layer:
-$\tilde{X}_{w}^{(i)}=W_{prj}\cdot X_{w}^{(i)}$ (3)
-where $W_{prj}\in\mathbb{R}^{w\times d}$ denotes the projection matrix of linear mapping, and $d$ represents the dimensionality of the hidden layer within the RNN structure.
-Undergoing linear projection transformation, the dimensionality of $\tilde{X}_{w}^{(i)}$ is defined as $\mathbb{R}^{C\times n\times d}$.
-However, this segmentation strategy has certain limitations.
-Once the segment length is chosen, the cut-off points are also determined.
-This may disrupt the continuity of information in time series analysis, thus limiting the model’s capabilities.
-The implicit segmentation method aims to break this limitation.
-Specifically, given a time series analysis vector $X^{(i)}\in\mathbb{R}^{L\times C}$, we first expand its last dimension and then swap the first and second dimensions to obtain $\tilde{X}^{(i)}\in\mathbb{R}^{C\times L\times 1}$ for subsequent processing.
-Then we utlize a linear layer to transform the last dimesntion of $\tilde{X}^{(i)}$ to get $\bar{X}(i)\in\mathbb{R}^{C\times L\times n}$ to represent the segmented matrix, subsequently, we interchange the second and third dimension, still denoted as $\bar{X}(i)$.
-Applying linear layer to the last dimension of $\bar{X}(i)$ yields $\bar{X}(i)\in\mathbb{R}^{C\times n\times d}$, which is consistent in dimension with the vector obtained via SegRNN methodology.
-Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation
-Notably, our proposed method does not explicitly define the segmentation length.
+Subsequently, we use linear mapping to transform the dimension $w$ into $d$, obtaining the embedded hidden state of the segmented vectors $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$.
+This segmentation method retains the sequence information while reducing the iteration count within the recurrent structure.
+However, this approach has some drawbacks.
+The simple truncation of data disrupts continuity, and there is no effective exchange of information between different segments, thus limiting the model’s capability.
+To address these issues, we propose an implicit segmentation strategy, as illustrated in Figure 2.
+This strategy initially decomposes the time series through linear mapping and subsequently transforms the information into the embedded hidden state $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$.
+Notably, after segmentation, each segment includes both the information of that segment and some additional information, ensuring more continuous information processing and thereby enhancing the model’s performance.
+Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation, we utilize a linear layers, from 1 to $n$, to decompose the time series.
+Subsequently, a second mapping is applied to compress the dimensions.
+Note that each segmented vector embedding can encompass information from other segments of the original time series.
+Furthermore, our proposed method does not explicitly define the segmentation length.
 Instead, it achieves data segmentation and dimensionality transformation through dual linear projections.
 The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information.
-It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
-4.3 Residual Structure for SegRNN
+It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe model but rather represents a general and straightforward technique for information enhancement.
+3.4 Residual Structure
 SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information.
@@
 SegRNN uses data segmentation to reduce the data’s flow through the RNN structure.
-Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbb{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
+Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
 $\mathbf{h}_{0}$ $=\mathbf{0}$ (6)
@@
 $\mathbf{h}_{n}$ $=f(\mathbf{h}_{n-1},\tilde{\mathbf{X}}_{w}^{(i)}[:,n,:];\mathbf{W})$
-where $\mathbf{h}_{t}\in\mathbb{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
+where $\mathbf{h}_{t}\in\mathbf{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
 Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$.
-However, the number of iterations still increases with the forecasting time steps.
-This unavoidably limits the model’s ability due to the loss of information during iterations.
-Therefore, we introduce a residual structure in the model, as shown in Figure 3.
-For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$.
-Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$.
-Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$.
-Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
-Figure 3: Residual block for SegRNN
-Introducing the residual structure allows some information to bypass the recurrent structure and be directly mapped to the encoder output.
+However, the number of iterations still increases with the forecasting time steps, which limits the model’s capability due to the loss of information during these iterations.
+To address this issue, we incorporate a residual structure into the model.
+For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbf{R}^{C\times L\times n}$, we apply a linear mapping to $\bar{\mathbf{X}}^{(i)}$, resulting in $\mathbf{h}^{\prime}_{n}$.
+Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$ to obtain the final output vector from the encoder.
+Introducing the residual structure allows some information to skip the recurrent structure and be directly mapped to the encoder output.
 This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.
@@
 Our experiments are structured as follows.
-In Section 5.1, we introduce the datasets used for long-term time series analysis.
-Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison.
-Section 5.3 presents and analyzes the performance metrics.
-In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model.
-Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model.
+We first introduce the datasets used for long-term time series analysis.
+Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison.
+Then we presents and analyzes the performance metrics.
+We further conducted ablation studies to investigate the effectiveness of the model.
+Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic.
 All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
 4.1 Datasets
-The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, and traffic, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings.
+The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings.
 Detailed dataset information is provided in Table 1.
 Table 1: Dataset Basic Attributes
-Dataset Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
-Time-Series 321 21 7 7 7 7
-Frequency 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
-Time-Points 26,304 52,696 17,420 17,420 69,680 69,680
+Data Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
+Chan. 321 21 7 7 7 7
+Freq. 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
+Points 26,304 52,696 17,420 17,420 69,680 69,680
 4.2 Experimental setup and Baselines
@@
 Note that all data, except for our method, originate from official code repositories and original papers.
-Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics.
-The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively.
+Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method and the iTransformer, each achieved top position in 7 and 6 metrics, respectively.
 This demonstrates the powerful capabilities of our model.
 Table 2: Performance metrics for Long-time series analysis.
-Datasets Models RNNs Transformers MLPs CNNs
+Datasets RNNs Transformers MLPs CNNs
 MSegRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 96 0.365 0.384 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
-192 0.415 0.413 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
+96 0.365 0.384 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
+ETTh1 192 0.415 0.413 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
 336 0.463 0.441 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
 720 0.468 0.460 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
-ETTh2 96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
-192 0.354 0.378 0.359 0.389 0.380 0.400 0.877 0.656 0.388 0.400 0.477 0.476 0.402 0.414
+96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
+ETTh2 192 0.354 0.378 0.359 0.389 0.380 0.400 0.877 0.656 0.388 0.400 0.477 0.476 0.402 0.414
 336 0.404 0.416 0.421 0.436 0.428 0.432 1.043 0.731 0.426 0.433 0.594 0.541 0.452 0.452
 720 0.407 0.428 0.432 0.455 0.427 0.445 1.104 0.763 0.431 0.446 0.831 0.657 0.462 0.468
-ETTm1 96 0.314 0.345 0.330 0.369 0.334 0.368 0.404 0.426 0.329 0.367 0.345 0.372 0.338 0.375
-192 0.361 0.374 0.369 0.392 0.377 0.391 0.450 0.451 0.367 0.385 0.380 0.389 0.374 0.387
+96 0.314 0.345 0.330 0.369 0.334 0.368 0.404 0.426 0.329 0.367 0.345 0.372 0.338 0.375
+ETTm1 192 0.361 0.374 0.369 0.392 0.377 0.391 0.450 0.451 0.367 0.385 0.380 0.389 0.374 0.387
 336 0.388 0.392 0.399 0.412 0.426 0.420 0.532 0.515 0.399 0.410 0.413 0.413 0.410 0.411
 720 0.448 0.430 0.454 0.443 0.491 0.459 0.666 0.589 0.454 0.439 0.474 0.453 0.478 0.450
-ETTm2 96 0.170 0.248 0.173 0.255 0.180 0.264 0.287 0.366 0.175 0.259 0.193 0.292 0.187 0.267
-192 0.233 0.292 0.237 0.298 0.250 0.309 0.414 0.492 0.241 0.302 0.284 0.362 0.249 0.309
+96 0.170 0.248 0.173 0.255 0.180 0.264 0.287 0.366 0.175 0.259 0.193 0.292 0.187 0.267
+ETTm2 192 0.233 0.292 0.237 0.298 0.250 0.309 0.414 0.492 0.241 0.302 0.284 0.362 0.249 0.309
 336 0.294 0.331 0.296 0.336 0.311 0.348 0.597 0.542 0.305 0.343 0.369 0.427 0.321 0.351
-720 0.391 0.389 0.389 0.407 0.412 0.407 1.730 1.042 0.402 0.400 0.554 0.522 0.408 0.403
-Electricity 96 0.149 0.242 0.151 0.245 0.148 0.240 0.219 0.314 0.181 0.270 0.197 0.282 0.168 0.272
-192 0.165 0.259 0.164 0.258 0.162 0.253 0.231 0.322 0.188 0.274 0.196 0.285 0.184 0.289
+720 0.391 0.389 0.389 0.407 0.412 0.407 1.024 0.679 0.394 0.390 0.387 0.370 0.400 0.388
+96 0.149 0.242 0.151 0.245 0.148 0.240 0.219 0.314 0.181 0.270 0.197 0.282 0.168 0.272
+Electricity 192 0.165 0.259 0.164 0.258 0.162 0.253 0.231 0.322 0.188 0.274 0.196 0.285 0.184 0.289
 336 0.181 0.275 0.180 0.277 0.178 0.269 0.246 0.337 0.204 0.293 0.209 0.301 0.198 0.300
 720 0.218 0.307 0.218 0.313 0.225 0.317 0.280 0.363 0.246 0.324 0.245 0.333 0.220 0.320
-Weather 96 0.155 0.193 0.165 0.227 0.174 0.214 0.158 0.230 0.177 0.210 0.196 0.255 0.172 0.220
-192 0.206 0.244 0.211 0.273 0.221 0.254 0.206 0.277 0.225 0.250 0.237 0.296 0.219 0.261
-336 0.260 0.288 0.270 0.318 0.278 0.296 0.272 0.335 0.278 0.290 0.283 0.335 0.280 0.306
-720 0.341 0.340 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
-Count 36 7 6 1 1 0 0
+96 0.147 0.187 0.165 0.227 0.174 0.214 0.158 0.230 0.177 0.210 0.196 0.255 0.172 0.220
+Weather 192 0.198 0.238 0.211 0.273 0.221 0.254 0.206 0.277 0.225 0.250 0.237 0.296 0.219 0.261
+336 0.256 0.282 0.270 0.318 0.278 0.296 0.272 0.335 0.278 0.290 0.283 0.335 0.280 0.306
+720 0.334 0.337 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
+Count 36 7 6 0 0 0 0
 We subsequently conducted experiments to observe how the model’s capability varies with changes in the look-back window.
@@
 We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model.
-Using the weather dataset as an example, we performed experiments excluding the Mamba structure and implicit segmentation structure respectivly.
+We performed experiments excluding the Mamba structure and implicit segmentation structure respectivly.
 For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure.
@@
 MSE MAE MSE MAE MSE MAE MSE MAE
-Weather 96 0.155 0.193 0.163 0.202 0.191 0.232 0.165 0.227
-192 0.206 0.244 0.207 0.242 0.243 0.277 0.211 0.273
-336 0.260 0.288 0.264 0.287 0.305 0.324 0.270 0.318
-720 0.341 0.335 0.350 0.342 0.380 0.377 0.357 0.376
-ETTh1 96 0.365 0.384 0.377 0.387 0.480 0.446 0.368 0.395
-192 0.415 0.413 0.423 0.416 0.520 0.472 0.408 0.419
+96 0.147 0.187 0.163 0.202 0.191 0.232 0.165 0.227
+Weather 192 0.199 0.238 0.207 0.242 0.243 0.277 0.211 0.273
+336 0.256 0.282 0.264 0.287 0.305 0.324 0.270 0.318
+720 0.332 0.337 0.350 0.342 0.380 0.377 0.357 0.376
+96 0.365 0.384 0.377 0.387 0.480 0.446 0.368 0.395
+ETTh1 192 0.415 0.413 0.423 0.416 0.520 0.472 0.408 0.419
 336 0.463 0.441 0.475 0.437 0.543 0.491 0.444 0.440
 720 0.468 0.460 0.476 0.454 0.549 0.515 0.446 0.457
-ETTh2 96 0.275 0.326 0.281 0.327 0.326 0.368 0.278 0.335
-192 0.354 0.378 0.368 0.385 0.418 0.420 0.359 0.389
+96 0.275 0.326 0.281 0.327 0.326 0.368 0.278 0.335
+ETTh2 192 0.354 0.378 0.368 0.385 0.418 0.420 0.359 0.389
 336 0.404 0.416 0.402 0.416 0.462 0.456 0.421 0.436
 720 0.407 0.428 0.421 0.435 0.467 0.472 0.432 0.455
-ETTm1 96 0.314 0.345 0.315 0.346 0.535 0.454 0.330 0.369
-192 0.361 0.374 0.362 0.376 0.556 0.469 0.369 0.392
+96 0.314 0.345 0.315 0.346 0.535 0.454 0.330 0.369
+ETTm1 192 0.361 0.374 0.362 0.376 0.556 0.469 0.369 0.392
 336 0.388 0.392 0.390 0.395 0.576 0.480 0.399 0.412
 720 0.448 0.430 0.446 0.427 0.631 0.509 0.454 0.443
-ETTm2 96 0.170 0.248 0.173 0.250 0.215 0.298 0.173 0.255
-192 0.233 0.292 0.235 0.293 0.283 0.339 0.237 0.298
+96 0.170 0.248 0.173 0.250 0.215 0.298 0.173 0.255
+ETTm2 192 0.233 0.292 0.235 0.293 0.283 0.339 0.237 0.298
 336 0.294 0.331 0.297 0.332 0.352 0.381 0.296 0.336
 720 0.391 0.389 0.392 0.388 0.463 0.441 0.389 0.407
-Average 0.335 0.355 0.341 0.357 0.425 0.411 0.340 0.372
-The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall.
-This is followed by the model that includes only implicit segmentation and residual structures.
-Interestingly, merely adding the Mamba structure does not significantly enhance model performance.
-This may be due to a synergistic effect between the two structures.
-Consequently, the ablation study confirms the superiority of our proposed structure.
+Average 0.334 0.355 0.341 0.357 0.425 0.411 0.340 0.372
 4.5 Model efficiency
-Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic.
+This section is dedicated to evaluating the model’s time and memory efficiency to substantiate its lightweight nature.
+The experiments are conducted on weather datasets because the other datasets utilize the Mamba architecture are implemented without convolutional layers and hardware optimization.
 The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments.
-Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Table 4.
-Table 4: Memory Usage of Two Methods
-Method Memory Usage (GB)
-96 192 336 720
-MSegRNN 0.98 1.17 1.56 2.48
-SegRNN 0.91 1.14 1.53 2.41
-Compare +7.7% +2.6% +2.0% +2.9%
-As shown in the table, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
+Using the Weather dataset as an example, the time and memory usage comparison with batch size 8 conducted on a single NVIDIA T4 GPU is shown in Figure 4.
+Figure 4: Comparsion of SegRNN and MSegRNN with time and space efficiency.
+As shown in the figure, compared to SegRNN, our model only increases memory usage and training time.
+However, in contrast to some transformer-based models, our method still holds significant spatiotemporal advantages.
 5 Conclusion
@@
 Appendix B Code Framework
-The implementation of our model framework is primarily derived from the official SegRNN code and the minimal implementation of Mamba.
+The implementation of our model framework is primarily derived from the official SegRNN code and the implementation of Mamba.
 This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output.
 We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
-The official SegRNN code is available at:https://github.com/lss-1138/SegRNN.
-The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal
+The official SegRNN source code is available at: https://github.com/lss-1138/SegRNN.
+The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal.
+The official implementation of Mamba can be found at:https://github.com/state-spaces/mamba.git
 Appendix C Detail Configuration
@@
 Table 4: Detail configuration of MSegRNN
-Datasets d_state seg_len d_model learning_rate dropout batch_size
-ETTh1 2 12 512 0.0003 0.1 256
-ETTh2 4 24 512 0.0015 0.5 256
-ETTm1 4 24 512 0.0010 0.1 256
-ETTm2 4 24 512 0.0010 0.1 256
-Weather 4 24 512 0.0007 0.1 64
-Electricity 2 48 512 0.0012 0.01 128
+Datasets d_state seg_len use_conv learning_rate dropout
+ETTh1 2 12 False 0.0003 0.1
+ETTh2 4 24 False 0.0015 0.5
+ETTm1 4 24 False 0.0010 0.1
+ETTm2 4 24 False 0.0010 0.1
+Weather 4 24 True 0.0007 0.1
+Electricity 2 48 False 0.0012 0.01
 We also employed a different learning rate adjustment strategy.
@@
 The result is shown in Table 5.
-Table 6: Perfromance of different models
+Table 5: Performance of different models
 Datasets Models MSegRNN TimeMachine Bi-Mamba4TS DTMamba
 MSE MAE MSE MAE MSE MAE MSE MAE
-Weather 96 0.155 0.193 0.164 0.208 0.164 0.212 0.171 0.218
-192 0.206 0.244 0.211 0.250 0.214 0.256 0.220 0.257
-336 0.260 0.288 0.256 0.290 0.269 0.296 0.274 0.296
-720 0.341 0.340 0.342 0.343 0.348 0.349 0.349 0.346
-ETTh1 96 0.365 0.384 0.364 0.387 0.376 0.403 0.386 0.399
-192 0.415 0.414 0.415 0.416 0.411 0.425 0.426 0.424
+96 0.147 0.187 0.164 0.208 0.164 0.212 0.171 0.218
+Weather 192 0.198 0.238 0.211 0.250 0.214 0.256 0.220 0.257
+336 0.256 0.282 0.256 0.290 0.269 0.296 0.274 0.296
+720 0.334 0.337 0.342 0.343 0.348 0.349 0.349 0.346
+96 0.365 0.384 0.364 0.387 0.376 0.403 0.386 0.399
+ETTh1 192 0.415 0.414 0.415 0.416 0.411 0.425 0.426 0.424
 336 0.463 0.441 0.429 0.421 0.455 0.445 0.480 0.450
 720 0.468 0.460 0.458 0.453 0.460 0.464 0.484 0.470
-ETTh2 96 0.275 0.326 0.275 0.334 0.289 0.341 0.290 0.340
-192 0.354 0.378 0.349 0.381 0.367 0.389 0.366 0.392
+96 0.275 0.326 0.275 0.334 0.289 0.341 0.290 0.340
+ETTh2 192 0.354 0.378 0.349 0.381 0.367 0.389 0.366 0.392
 336 0.404 0.416 0.340 0.381 0.410 0.424 0.380 0.409
 720 0.407 0.428 0.411 0.433 0.421 0.439 0.416 0.437
-ETTm1 96 0.314 0.345 0.317 0.355 0.312 0.354 0.325 0.360
-192 0.361 0.374 0.357 0.378 0.358 0.383 0.375 0.386
+96 0.314 0.345 0.317 0.355 0.312 0.354 0.325 0.360
+ETTm1 192 0.361 0.374 0.357 0.378 0.358 0.383 0.375 0.386
 336 0.388 0.392 0.379 0.399 0.388 0.404 0.396 0.405
 720 0.448 0.430 0.445 0.436 0.444 0.433 0.454 0.442
-ETTm2 96 0.170 0.248 0.175 0.256 0.174 0.259 0.177 0.259
-192 0.233 0.292 0.239 0.299 0.240 0.304 0.240 0.300
+96 0.170 0.248 0.175 0.256 0.174 0.259 0.177 0.259
+ETTm2 192 0.233 0.292 0.239 0.299 0.240 0.304 0.240 0.300
 336 0.294 0.331 0.287 0.332 0.303 0.345 0.310 0.345
 720 0.391 0.389 0.371 0.385 0.402 0.407 0.395 0.394
-Electricity 96 0.149 0.242 0.142 0.236 0.142 0.238 0.166 0.256
-192 0.165 0.259 0.158 0.250 0.157 0.253 0.178 0.268
+96 0.149 0.242 0.142 0.236 0.142 0.238 0.166 0.256
+Electricity 192 0.165 0.259 0.158 0.250 0.157 0.253 0.178 0.268
 336 0.181 0.275 0.172 0.268 0.172 0.271 0.197 0.289
@@
 Additionally, the Bi-Mamba4TS method also demonstrates a clear advantage on the Electricity dataset.
+Appendix E Prediction Result
+We subsequently observed the partial prediction results of the model using the weather dataset, with a look-back window of 96 and a prediction horizon of 336. The results are shown in Figure 5:
+Figure 5: Prediction result of different method.
+The four methods are Dlinear, MSegRNN, SegRNN, and PatchTST in sequence.
+Intuitively, our method also has the best fitting performance, which is consistent with the fact that MSegRNN has the minimum loss among the models.
+Appendix F Ablation study on using convolution layers
+In the main text’s model structure, we removed the convolutional layer from the Mamba structure due to performance issues.
+We hypothesized that local features might have been redundantly captured.
+To investigate the impact of the convolutional layer on the model performence, we conducted experiments.
+The results are shown in Table 6.
+Table 6: Ablation study on convlution layer
+Datasets Models WithoutConv WithConv
+MSE MAE MSE MAE
+96 0.365 0.384 0.412 0.426
+ETTh1 192 0.415 0.414 0.444 0.443
+336 0.463 0.441 0.477 0.460
+720 0.468 0.460 0.502 0.492
+96 0.275 0.326 0.282 0.338
+ETTh2 192 0.354 0.378 0.372 0.396
+336 0.404 0.416 0.406 0.424
+720 0.407 0.428 0.430 0.456
+96 0.314 0.345 0.323 0.355
+ETTm1 192 0.361 0.374 0.373 0.384
+336 0.388 0.392 0.397 0.401
+720 0.448 0.430 0.452 0.432
+96 0.170 0.248 0.180 0.264
+Ettm2 192 0.233 0.292 0.240 0.305
+336 0.294 0.331 0.299 0.343
+720 0.391 0.389 0.396 0.399
+It can be observed that except for the Weather dataset in our experiments, the convolutional layer may interfere with the model’s predictions in most scenarios, thereby limiting the model’s capability.

### v4

@@
-MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting
+ISMRNN: An Implicitly Segmented RNN Method with Mamba for Long-Term Time Series Forecasting
 Abstract
@@
 Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure.
-To address these issues, we propose the MSegRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase.
+To address these issues, we propose the ISMRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase.
 Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure.
@@
 In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy.
-Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
+Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed ISMRNN, the main contributions of ISMRNN can be summarized as follows:
   • Introducing the Mamba structure to effectively preprocessing the information.
@@
 Additionally, segmentation methods still face the challenge of cumulative information processing within the recurrent structure.
-In this paper, we propose the MSegRNN method to overcome these limitations through three strategies, as illustrated in Figure 1.
+In this paper, we propose the ISMRNN method to overcome these limitations through three strategies, as illustrated in Figure 1.
 We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage.
@@
 We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
-Figure 1: Overall structure of MSegRNN model
+Figure 1: Overall structure of ISMRNN model
 3.2 Mamba structure
@@
 Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic.
-All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
+All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/ISMRNN.git.
 4.1 Datasets
@@
 Datasets RNNs Transformers MLPs CNNs
-MSegRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
+ISMRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 The evaluation was performed on the ETTh2 dataset, with the MSE and MAE losses of the method shown in the Figure 3:
-Figure 3: Comparsion of SegRNN and MSegRNN with vary look-back window length
+Figure 3: Comparsion of SegRNN and ISMRNN with vary look-back window length
 Our model exhibited smaller losses across various look-back windows, demonstrating the superiority of our method.
@@
 Using the Weather dataset as an example, the time and memory usage comparison with batch size 8 conducted on a single NVIDIA T4 GPU is shown in Figure 4.
-Figure 4: Comparsion of SegRNN and MSegRNN with time and space efficiency.
+Figure 4: Comparsion of SegRNN and ISMRNN with time and space efficiency.
 As shown in the figure, compared to SegRNN, our model only increases memory usage and training time.
@@
 5 Conclusion
-In this work, we present the MSegRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework.
+In this work, we present the ISMRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework.
 These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model.
@@
 The results demonstrate that our model yields significant imporvements in some real-world scenarios.
-To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of MSegRNN against SegRNN on datasets of varying scales.
+To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of ISMRNN against SegRNN on datasets of varying scales.
 The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
-Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the MSegRNN method excels with corresponding empirical research.
+Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the ISMRNN method excels with corresponding empirical research.
 Appendix
@@
 The parameter settings and results for other methods are taken from their original papers or official code repositories.
-Table 4: Detail configuration of MSegRNN
+Table 4: Detail configuration of ISMRNN
 Datasets d_state seg_len use_conv learning_rate dropout
@@
 Table 5: Performance of different models
-Datasets Models MSegRNN TimeMachine Bi-Mamba4TS DTMamba
+Datasets Models ISMRNN TimeMachine Bi-Mamba4TS DTMamba
 MSE MAE MSE MAE MSE MAE MSE MAE
@@
 Figure 5: Prediction result of different method.
-The four methods are Dlinear, MSegRNN, SegRNN, and PatchTST in sequence.
-Intuitively, our method also has the best fitting performance, which is consistent with the fact that MSegRNN has the minimum loss among the models.
+The four methods are Dlinear, ISMRNN, SegRNN, and PatchTST in sequence.
+Intuitively, our method also has the best fitting performance, which is consistent with the fact that ISMRNN has the minimum loss among the models.
 Appendix F Ablation study on using convolution layers